Show that y=x+tanx satisfies the differential equation cos2x d2y /dx2 -2y+2x=0


1
Expert's answer
2021-01-07T15:33:14-0500

Given, "y=x+\tan x ...(i)"

On differentiating both sides w.r.t "x" ,

"\frac{dy}{dx}=1+\sec ^2x"

Again, differentiating both sides w.r.t "x" ,

"\frac{d^2y}{dx^2}=0+2\sec x \frac{d (\sec x)}{dx}=2\sec x \sec x\tan x
\\\Rightarrow \frac{d^2y}{dx^2}=2\sec^2 x \tan x ...(ii)"

Now, differential equation is: "\cos ^2x \frac{d^2y}{dx^2}-2y+2x=0"

Put values from equation (i) and (ii) into LHS of this, we get:

"=\cos ^2x (2\sec^2 x \tan x)-2(x+\tan x)+2x"

"=\cos ^2x (2 \times\frac{1}{\cos ^2x} \tan x)-2x-2\tan x+2x \ [\because \ \sec x=\frac{1}{\cos x}]
\\ =2 \tan x-2x-2\tan x+2x
\\ =0+0
\\ =0
\\ =RHS"

Hence, proved.